Ex 6.1 Class 8 Maths Question 9.
How many numbers lie between squares of the following numbers?
(i) 12 and 13
(ii) 25 and 26
(iii) 99 and 100.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(i) We know that numbers between n2 and (n + 1)2 = 2n Numbers between 122 and 132 = (2n) = 2 × 12 = 24 (ii) Numbers between 252 and 262 = 2 × 25 = 50 (∵ n = 25) (iii) Numbers between 992 and 1002 = 2 × 99 = 198 (∵ n = 99)